A Mercury in glass thermometer reads -200C at the ice point and 1000C at the steam point, Calculate the Celcius Temperature corresponding to 70° on the thermometer


1
Expert's answer
2021-04-28T13:03:34-0400

Find how the temperature in the strange thermometer depends on the normal Celsius scale. Assume x is Celsius scale and y is strange scale:


"y-y_1=m(x-x_1),\\
y-(-200)=m(x-0),\\
m=\frac{1000-(-200)}{100-0}=12.\\\space\\
y=12x-200."

So, 70°C will be


"y=12\u00b770-200=4C"

on the strange thermometer.